# Energy-Aware Lane Planning for Connected Electric Vehicles in Urban Traffic: Design and Vehicle-in-the-Loop Validation

### v2

@@ I Introduction
Connected and Automated Vehicles (CAVs) can improve safety, traffic flow, and energy efficiency by coordinating with traffic signals and nearby vehicles through vehicle-to-infrastructure (V2I) and vehicle-to-vehicle (V2V) communication [1]. Eco-driving strategies leveraging Signal Phase and Timing (SPaT) data from connected traffic lights have demonstrated 11–16% energy savings [2, 3], but these methods typically assume fixed-lane driving and optimize only longitudinal motion [4, 5, 6]. In practice, lane changes also influence efficiency by avoiding slower vehicles or aligning with favorable traffic signals [5]. Yet most eco-driving studies overlook this factor, limiting their applicability to realistic multi-lane scenarios. We extend eco-driving to include lane-change decisions for energy savings for CAVs. We focus on how connectivity (V2I/V2V) can help an energy-aware lane selection strategy, extending conventional eco-driving to a multi-lane context.
@@ II-A Eco-Driving with V2I Connectivity
A large body of research has explored energy-efficient driving using V2I communication, particularly along signalized corridors. Optimization-based eco-driving controllers have been proposed to leverage SPaT data over corridors with multiple traffic lights. For example, in [7], the authors apply Pontryagin’s Minimum Principle to develop an eco-driving controller, demonstrating fuel savings of up to 36% compared to a human-modeled driver through VIL experiments. In [8], the authors design a real-time eco-driving controller formulated as a receding-horizon optimal control problem and demonstrate its effectiveness through real-world road testing, achieving a 31% energy reduction at the cost of increased travel time. These works demonstrate that CAVs can significantly improve efficiency through longitudinal speed planning, when leveraging infrastructure information. However, these studies assume a fixed lane. The controllers react to surrounding traffics and red lights, but do not consider proactive lane changes to avoid deceleration.
@@ II-B Lane Changes in Eco-Driving
Only recently have researchers begun to incorporate lateral maneuvers into eco-driving. In [9], the authors note that most studies emphasize longitudinal control and neglect lane changes, which can reduce efficiency when following slower vehicles. To address this, they propose a graph-search strategy that selects lane sequences based on predicted traffic, though the effect of traffic lights is not considered. Similarly, [10] designs an overtaking strategy in which an energy-optimal speed trajectory is first computed, and a lane change is triggered if a slower vehicle disrupts that plan. These studies indicate that lane-changing can enhance longitudinal eco-driving, yet their validation remains limited to simulation. Few works have demonstrated lane-changing eco-driving in real vehicles or high-fidelity VIL platforms, aside from our previous work [11].
@@ IV Problem Setup
Figure 1: Vehicle-in-the-loop control architecture [11]
In this section, we formulate the problem of energy-efficient lane selection for connected, automated urban driving. We adopt the hierarchical control architecture introduced in [11] and focus on designing an energy-efficient lane selector for battery electric vehicles (BEVs) in urban environments. The architecture, illustrated in Fig. 1, and the corresponding vehicle-in-the-loop (VIL) setup are detailed in [11]. We assume that a route from point A to point B is given and that the traffic light cycles are deterministic.
@@ IV-B Energy-efficient Lane Selection
IV-B1 Data-driven Energy Consumption Model [6]
We approximate the energy consumption of the battery electric vehicle (BEV) given a longitudinal trajectory. Specifically, we assume that the energy consumption is a function of the longitudinal displacement along the centerline and the longitudinal acceleration of the vehicle. For more details, refer to Sec. II.C. in [6]. By recording the vehicle’s state-of-charge (SOC) and longitudinal motion, we fit a quadratic regression model to estimate the energy consumption rate $\Delta E$, denoted as $\mathcal{E}(v,a)$, where $v$ and $a$ are the vehicle’s longitudinal velocity and acceleration. The quadratic regression model is parameterized as
@@ IV-B1 Data-driven Energy Consumption Model [6]
The resulting regression and the corresponding measured energy consumption data are shown in Fig. 3. The mean error is approximately 1% [12]. The observed discrepancy between the model and the measured data is attributed to unmodeled non-linear effects such as aerodynamic drag, rolling resistance, and variations in motor and battery efficiency. Note that this approximate energy consumption model is used to reason about energy consumption for lane selection, but not used to measure the actual energy consumption of the vehicle. We directly read and record the SOC from the vehicle’s CAN signals.
@@ V-A2 Graph Search Terminal Cost Approximation
where $v(V)$ represents the passing velocity value of the node and $\ell$ denotes the $\ell$-th traffic light that the node $V_{j}$ represents. $tl^{\ell,p}$ at the time the ego vehicle reaches the traffic light is computed based on the travel time between nodes assuming the vehicle travels at a constant velocity between traffic lights. $\mathbb{1}(\cdot)$ is the Heaviside step function. The connections from nodes in $\mathcal{V}_{N_{TL}}$ to the terminal node $V_{-1}$ are added to penalize the velocity deviation from the velocity at the point of passing the first traffic light. Alternatively, the data-driven energy model in (5) can be used to approximate the energy consumption resulting from changes in node velocities, assuming constant acceleration. However, for simplicity, this work adopts the kinetic energy model to compute edge costs. Note that the graph (9) satisfies additive property meaning $c(V_{a},V_{c})=c(V_{a},V_{b})+c(V_{b},V_{c})$ for any node connections $V_{a}\rightarrow V_{b}$ and $V_{b}\rightarrow V_{c}$. Additivity allows us to exploit dynamic programming and optimal substructure, formulating the problem as a path-finding problem and applying Dijkstra’s algorithm [13] to compute $\min c(V_{1},V_{-1})$ given $v(V_{1})$. We define the minimum sum of the edge cost from the starting node to the terminal node as
$J_{graph}(v(V_{1}))=\min_{\{v_{i}\in\mathcal{V}_{i}\}_{i=2}^{N_{TL}}}c(V_{1},V_{-1}),$ (11)
@@ V-A3 Constraints
Collision avoidance constraints with respect to the NPC vehicles are also considered, and their set representation is denoted as $\mathcal{F}^{LK}(s^{\text{front}},v^{\text{front}})$, where $s^{\text{front}}$ and $v^{\text{front}}$ denote the states of the vehicle immediately preceding the ego vehicle in the same lane. We assume that the NPC vehicles maintain constant velocity during the prediction horizon $N$. For more details, refer to [11].
@@ V-A4 Optimal Control Problem
We solve the following constrained finite-time optimal control problem for every $\mathcal{C}_{\text{pass}}$ candidates in the set (3) and collect the optimal cost $J^{\star}_{\text{ftocp}}(x(t),u(t))$:
@@
$\quad(x_{k|t},u_{k|t})\in\mathcal{F}^{LK}(s^{\text{front}}_{k|t},v^{\text{front}}_{k|t})$ (19d)
$\quad x_{t|t}=x(t),u_{t|t}=u(t),\leavevmode\nobreak\ \forall k\in\mathbb{I}_{t}^{t+N-1}$ (19e)
$\quad x_{k|t}\in\mathcal{P}(tl^{\text{next},s}),\leavevmode\nobreak\ \forall k\in\mathbb{I}_{[t+\underline{t}_{\text{pass}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}^{[t+\bar{t}_{\text{pass}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}$ (19f)
$\quad x_{k|t}\in\mathcal{S}(tl^{\text{next},s}),\leavevmode\nobreak\ \forall k\in\mathbb{I}_{[t+\underline{t}_{\text{stop}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}^{[t+\bar{t}_{\text{stop}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}$ (19g)
where $x_{k|t}$ denotes the decision variable at time step $k$ predicted at time $t$ and $N$ is the prediction horizon. In this work, we use $N=140$ to achieve 1 $Hz$ lane selection planning with the available hardware. The $\mathbb{I}_{k_{1}}^{k_{2}}$ notation denotes the set $\{k_{1},k_{1}+1,\dots,k_{2}\}$. The optimal control problem has three components for the cost function. The first component is the energy minimization cost which uses the data-driven energy consumption model: $\mathcal{E}$, which is quadratic and convex w.r.t. the decision variables. The second component $J^{LK}_{\text{smooth}}(\{x_{k|t}\}^{t+N-1}_{t})$ is the stage cost added for smoothness of the longitudinal trajectory and passenger comfort [11]. After the optimal control problem is formulated in CasADi [14] and solved using IPOPT [15], the passing velocity of the first traffic light $v^{\star}_{\text{pass}}$ is extracted from the optimal solution, which is used to initialize the graph (9). After computing (11), we define
$J^{\star}($ $\mathcal{C}_{\text{pass}})=J_{\text{graph}}(v^{\star}_{\text{pass}})+J^{\star}_{\text{ftocp}}(x(t),u(t)).$ (20)
@@
The lane selector also provides the reference velocity profile $v_{\text{ref}}$, which represents the energy-efficient speed trajectory corresponding to the selected behavior decision $\mathcal{C}^{\text{eco}}_{\text{pass}}$. The lower-level trajectory planner then computes a motion trajectory that closely tracks this reference velocity while respecting traffic laws and ensuring safety with respect to surrounding vehicles, using a finer time discretization to guarantee dynamic feasibility and reactivity [11].
@@ VI-1 Experimental Setup
We evaluate the proposed approach using a Vehicle-in-the-Loop (VIL) system that integrates a Hyundai IONIQ 5 with CARLA simulation for realistic traffic and SPaT signals. Fig. 5 shows the virtual CARLA simulator map, the satellite image of the physical testing site, and the test vehicle (IONIQ 5). This setup enables robust evaluation of planning and control algorithms in a safe environment while using real energy measurements on the physical vehicle. See [11] for hardware details.

### v3

@@ I Introduction
Connected and Automated Vehicles (CAVs) can improve safety, traffic flow, and energy efficiency by coordinating with traffic signals and nearby vehicles through vehicle-to-infrastructure (V2I) and vehicle-to-vehicle (V2V) communication [GUANETTI201818]. Eco-driving strategies leveraging Signal Phase and Timing (SPaT) data from connected traffic lights have demonstrated 11–16% energy savings [Levin2023, Rousseau], but these methods typically assume fixed-lane driving and optimize only longitudinal motion [farias, sangjae, joa_ioniq5]. In practice, lane changes also influence efficiency by avoiding slower vehicles or aligning with favorable traffic signals [sangjae]. Yet most eco-driving studies overlook this factor, limiting their applicability to realistic multi-lane scenarios. We extend eco-driving to include lane-change decisions for energy savings for CAVs. We focus on how connectivity (V2I/V2V) can help an energy-aware lane selection strategy, extending conventional eco-driving to a multi-lane context.
@@ II-A Eco-Driving with V2I Connectivity
A large body of research has explored energy-efficient driving using V2I communication, particularly along signalized corridors. Optimization-based eco-driving controllers have been proposed to leverage SPaT data over corridors with multiple traffic lights. For example, in [ard2023VILCAV], the authors apply Pontryagin’s Minimum Principle to develop an eco-driving controller, demonstrating fuel savings of up to 36% compared to a human-modeled driver through VIL experiments. In [bae2022ecological], the authors design a real-time eco-driving controller formulated as a receding-horizon optimal control problem and demonstrate its effectiveness through real-world road testing, achieving a 31% energy reduction at the cost of increased travel time. These works demonstrate that CAVs can significantly improve efficiency through longitudinal speed planning, when leveraging infrastructure information. However, these studies assume a fixed lane. The controllers react to surrounding traffics and red lights, but do not consider proactive lane changes to avoid deceleration.
@@ II-B Lane Changes in Eco-Driving
Only recently have researchers begun to incorporate lateral maneuvers into eco-driving. In [dong2024lc], the authors note that most studies emphasize longitudinal control and neglect lane changes, which can reduce efficiency when following slower vehicles. To address this, they propose a graph-search strategy that selects lane sequences based on predicted traffic, though the effect of traffic lights is not considered. Similarly, [yang2023lc] designs an overtaking strategy in which an energy-optimal speed trajectory is first computed, and a lane change is triggered if a slower vehicle disrupts that plan. These studies indicate that lane-changing can enhance longitudinal eco-driving, yet their validation remains limited to simulation. Few works have demonstrated lane-changing eco-driving in real vehicles or high-fidelity VIL platforms, aside from our previous work [arpae-lc].
@@ IV Problem Setup
Figure 1: Vehicle-in-the-loop control architecture [arpae-lc]
In this section, we formulate the problem of energy-efficient lane selection for connected, automated urban driving. We adopt the hierarchical control architecture introduced in [arpae-lc] and focus on designing an energy-efficient lane selector for battery electric vehicles (BEVs) in urban environments. The architecture, illustrated in Fig. 1, and the corresponding vehicle-in-the-loop (VIL) setup are detailed in [arpae-lc]. We assume that a route from point A to point B is given and that the traffic light cycles are deterministic.
@@ IV-B Energy-efficient Lane Selection
IV-B1 Data-driven Energy Consumption Model [joa_ioniq5]
We approximate the energy consumption of the battery electric vehicle (BEV) given a longitudinal trajectory. Specifically, we assume that the energy consumption is a function of the longitudinal displacement along the centerline and the longitudinal acceleration of the vehicle. For more details, refer to Sec. II.C. in [joa_ioniq5]. By recording the vehicle’s state-of-charge (SOC) and longitudinal motion, we fit a quadratic regression model to estimate the energy consumption rate $\Delta E$, denoted as $\mathcal{E}(v,a)$, where $v$ and $a$ are the vehicle’s longitudinal velocity and acceleration. The quadratic regression model is parameterized as
@@ IV-B1 Data-driven Energy Consumption Model [joa_ioniq5]
The resulting regression and the corresponding measured energy consumption data are shown in Fig. 3. The mean error is approximately 1% [joa2024ioniq5journal]. The observed discrepancy between the model and the measured data is attributed to unmodeled non-linear effects such as aerodynamic drag, rolling resistance, and variations in motor and battery efficiency. Note that this approximate energy consumption model is used to reason about energy consumption for lane selection, but not used to measure the actual energy consumption of the vehicle. We directly read and record the SOC from the vehicle’s CAN signals.
@@ V-A2 Graph Search Terminal Cost Approximation
where $v(V)$ represents the passing velocity value of the node and $\ell$ denotes the $\ell$-th traffic light that the node $V_{j}$ represents. $tl^{\ell,p}$ at the time the ego vehicle reaches the traffic light is computed based on the travel time between nodes assuming the vehicle travels at a constant velocity between traffic lights. $\mathbb{1}(\cdot)$ is the Heaviside step function. The connections from nodes in $\mathcal{V}_{N_{TL}}$ to the terminal node $V_{-1}$ are added to penalize the velocity deviation from the velocity at the point of passing the first traffic light. Alternatively, the data-driven energy model in (5) can be used to approximate the energy consumption resulting from changes in node velocities, assuming constant acceleration. However, for simplicity, this work adopts the kinetic energy model to compute edge costs. Note that the graph (9) satisfies additive property meaning $c(V_{a},V_{c})=c(V_{a},V_{b})+c(V_{b},V_{c})$ for any node connections $V_{a}\rightarrow V_{b}$ and $V_{b}\rightarrow V_{c}$. Additivity allows us to exploit dynamic programming and optimal substructure, formulating the problem as a path-finding problem and applying Dijkstra’s algorithm [dijkstra1959note] to compute $\min c(V_{1},V_{-1})$ given $v(V_{1})$. We define the minimum sum of the edge cost from the starting node to the terminal node as
$J^{\star}_{graph}(v(V_{1}))=\min_{\{V_{i}\in\mathcal{V}_{i}\}_{i=2}^{N_{TL}}}c(V_{1},V_{-1}),$ (11)
@@ V-A3 Constraints
Collision avoidance constraints with respect to the NPC vehicles are also considered, and their set representation is denoted as $\mathcal{F}^{LK}(s^{\text{front}},v^{\text{front}})$, where $s^{\text{front}}$ and $v^{\text{front}}$ denote the states of the vehicle immediately preceding the ego vehicle in the same lane. We assume that the NPC vehicles maintain constant velocity during the prediction horizon $N$. For more details, refer to [arpae-lc].
@@ V-A4 Optimal Control Problem
We solve the following constrained finite-time optimal control problem for every $\mathcal{C}_{\text{pass}}$ candidates in the set (3) and collect the optimal cost $J^{\star}_{\text{ftocp}}(x(t),u(t),\mathcal{C}_{\text{pass}})$:
@@
$\quad(x_{k|t},u_{k|t})\in\mathcal{F}^{LK}(s^{\text{front}}_{k|t},v^{\text{front}}_{k|t}),$ (19d)
$\quad x_{t|t}=x(t),\penalty 10000\ \forall k\in\mathbb{I}_{t}^{t+N-1}$ (19e)
$\quad x_{k|t}\in\mathcal{P}(tl^{\text{next},s}),\penalty 10000\ \forall k\in\mathbb{I}_{[t+\underline{t}_{\text{pass}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}^{[t+\bar{t}_{\text{pass}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}$ (19f)
$\quad x_{k|t}\in\mathcal{S}(tl^{\text{next},s}),\penalty 10000\ \forall k\in\mathbb{I}_{[t+\underline{t}_{\text{stop}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}^{[t+\bar{t}_{\text{stop}}(\mathcal{C}_{\text{pass}},tl^{\text{next},p})]/\Delta t}$ (19g)
where $x_{k|t}$ denotes the decision variable at time step $k$ predicted at time $t$ and $N$ is the prediction horizon. In this work, we use $N=140$ to achieve 1 $Hz$ lane selection planning with the available hardware. The $\mathbb{I}_{k_{1}}^{k_{2}}$ notation denotes the set $\{k_{1},k_{1}+1,\dots,k_{2}\}$. The optimal control problem has three components for the cost function. The first component is the energy minimization cost which uses the data-driven energy consumption model: $\mathcal{E}$, which is quadratic and convex w.r.t. the decision variables. The second component $J^{LK}_{\text{smooth}}(\{x_{k|t}\}^{t+N-1}_{t})$ is the stage cost added for smoothness of the longitudinal trajectory and passenger comfort [arpae-lc]. After the optimal control problem is formulated in CasADi [casadi] and solved using IPOPT [ipopt], the passing velocity of the first traffic light $v^{\star}_{\text{pass}}$ is extracted from the optimal solution, which is used to initialize the graph (9). After computing (11), we define
$J^{\star}($ $\mathcal{C}_{\text{pass}})=J_{\text{graph}}^{\star}(v^{\star}_{\text{pass}})+J^{\star}_{\text{ftocp}}(x(t),u(t),\mathcal{C}_{\text{pass}}).$ (20)
@@
The lane selector also provides the reference velocity profile $v_{\text{ref}}$, which represents the energy-efficient speed trajectory corresponding to the selected behavior decision $\mathcal{C}^{\text{eco}}_{\text{pass}}$. The lower-level trajectory planner then computes a motion trajectory that closely tracks this reference velocity while respecting traffic laws and ensuring safety with respect to surrounding vehicles, using a finer time discretization to guarantee dynamic feasibility and reactivity [arpae-lc].
@@ VI-1 Experimental Setup
We evaluate the proposed approach using a Vehicle-in-the-Loop (VIL) system that integrates a Hyundai IONIQ 5 with CARLA simulation for realistic traffic and SPaT signals. Fig. 5 shows the virtual CARLA simulator map, the satellite image of the physical testing site, and the test vehicle (IONIQ 5). This setup enables robust evaluation of planning and control algorithms in a safe environment while using real energy measurements on the physical vehicle. See [arpae-lc] for hardware details.